OpenStudy (anonymous):

how would you solve -1=1000-x all over x

OpenStudy (tkhunny):

This: \(-1 = \dfrac{1000-x}{x}\)

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

There are a few ways to go about it. 1) See in your mind that no matter what else happents \(x \ne 0\). 2) Rewrite a little... \(-1 = \dfrac{1000}{x} - 1\) Do you see where all that came from?

OpenStudy (anonymous):

it looks like you took the -x/x and made it -1. the x under the 1000 stays because the whole thing can be rewritten as 1000/x and -x/x

OpenStudy (anonymous):

let me chevk to make sure that's the correct p

OpenStudy (tkhunny):

Perfect. And then -x/x = -1. We're not done, yet.

OpenStudy (anonymous):

problem because if you had one to both sides you're left with 0=1000/x

OpenStudy (tkhunny):

Excellent. Why do you think that is a problem?

OpenStudy (anonymous):

there would be no solution because it will become 0=1000

OpenStudy (anonymous):

alright so the correct problem was -1= - 1000-1 over x

OpenStudy (tkhunny):

I think you have it. No solution. Done. That's not a problem. That's a solution. However, working the right problem will help. Let's see what you get,

OpenStudy (anonymous):

thank you so much for clearing that up. that stumped me for a while. Are you good with optimizations?

OpenStudy (tkhunny):

Not if I don't see the problem statement. Post a new thread.

OpenStudy (anonymous):

ok before i do that. still looking at this one. even with the first term being a negative I dont see it affecting the x. should I factor out the negative?

OpenStudy (anonymous):

I keep going back because I know that x = 500

OpenStudy (tkhunny):

It is of no consequence. Do or don't. Just do it right. -1= - 1000-1 over x Is this \(-1 = -1000 - \dfrac{1}{x}\)

OpenStudy (anonymous):

okay so it isnt \[-1= \frac{ -1000 }{ x } - \frac{ 1 }{ x } but -1 = -1000 -\frac{ 1 }{ x }\]

OpenStudy (tkhunny):

Okay, so then 999 = -1/x and x = -1/999 Done.

OpenStudy (anonymous):

no thats not right either. I'm looking at my notes and it's showing 500 to be the answer. Analytically that makes sense, but I can't get the math behind it

OpenStudy (anonymous):

thsnk you

OpenStudy (tkhunny):

It always helps to start with the right question. I have no confidence that we are doing that.